DogGirl27:

Rounded to the nearest hundredth, what is the positive solution to the quadratic equation 0=2x2+3x-8? Quadratic formula: x =StartFraction negative b plus or minus StartRoot b squared minus 4 a c EndRoot Over 2 a EndFraction 1.39 2.00 2.89 3.50

snowflake0531:

\[Quadratic ~equation~: ~~ \frac{-b \pm \sqrt{b^2-4ac}}{2a}\] \[a=2\\b=3\\c=-9\]

snowflake0531:

*-8

DogGirl27:

i dont understand

DogGirl27:

so what is the answer?

snowflake0531:

\[\frac{-3 \pm \sqrt{3^2-4(2)(-9)}}{2(2)}\] and simplify that

Alex1513:

The solution is 1.39 rounded up to the nearest hundredth

DogGirl27:

Thank you so much

Alex1513:

np

GoodFellow:

@doggirl27 wrote:
Thank you so much
You need just the answer, don't you?